# Incompressible Euler equations in 3D bounded domains in a critical space

### 30 Mar 2026 v1

@@ Abstract.
We consider the 3D incompressible Euler equations in bounded domains $\Omega$ with smooth boundary $\partial\Omega$. Based on [IMT-2019], we define the Besov space $B^{s}_{p,q}(A)$ by means of the Stokes operator $A$ with the Neumann boundary condition on $\partial\Omega$, and prove unique local existence theorem of strong solution for the initial data in the critical Besov space $B^{\frac{5}{2}}_{2,1}(A)$. Our proof relies on the method of vanishing viscosity. The commutator estimate plays an essential role for derivation of energy bounds which hold uniformly with respect to viscosity constants.
@@ 1. Introduction
In the present paper, we prove the local well-posedness in the critical Besov space $B^{\frac{5}{2}}_{2,1}(\Omega)$. Based on [IwKo-preprint] and [IMT-2019], we first define the Besov space $B^{s}_{p,q}(A)$ in terms of the Stokes operator $A$ in $L^{2}_{\sigma}(\Omega)$ with the Neumann boundary condition on $\partial\Omega$. Our method relies on the invisid limit ${\lim_{\varepsilon\to+0}u_{\varepsilon}}$ of a family $\{u_{\varepsilon}\}_{\varepsilon>0}$ of solutions to the following Navier-Stokes equations;
@@
It is essential to establish uniform bounds of $\{u_{\varepsilon}\}_{\varepsilon>0}$ in the Chemin-Lernar space $\tilde{L}^{\infty}(0,T;B^{\frac{5}{2}}_{2,1}(A))$. The commutator type estimate such as Kato-Ponce [KaPo-1988] plays a crucial role for such bounds. Since we make use of the Stokes operator $A$ with the Neumann boundary condition like (1.2), the boundary layer does not appear in the invisid limit.
@@
$\left\|u-\sum_{j\leq J}\phi_{j}(\sqrt{A})u\right\|_{\dot{B}^{\frac{5}{2}}_{2,1}(A)}\leq C2^{-\frac{5}{2}J}$
@@
This paper is organized as follows. In Section 2, we provide several preliminary lemmas and propositions required to prove Theorem 1.1. Section 3 is devoted to constructing solutions for the equation regularized by $\varepsilon\Delta$. In Section 4, we establish the uniform boundedness of the solutions obtained in Section 3 with respect to the parameter $\varepsilon>0$. Finally, in Section 5, we provide the proof of Theorem 1.1 by taking the limit as $\varepsilon\to 0$. Throughout the following sections, we suppose that $A_{2}$ is strictly positive.
@@ 2. Preliminary
In this section, we collect several preliminary definitions and results. In Subsection 2.1, we fundamental properties of the Stokes operator $A$ and the associated Besov spaces $B^{s}_{p,q}(A)$. Subsection 2.2 is devoted to providing several propositions and lemmas required for the proof of Theorem 1.1, including Sobolev-type embeddings, derivative estimates, and various properties of derivatives involving spectral restriction operators.
@@ 2.1. The Stokes operator and Besov spaces
We introduce the projection $\mathbb{P}$ onto the space of divergence-free vector fields and the Stokes operator on $\mathcal{X}_{\sigma}^{\prime}$. While the operator $A$ and its resolvent act on divergence-free fields, $B$ and its resolvent act on a more general class of vector fields.
Definition. (The projection $\mathbb{P}$, the Stokes operator $A$ and the Laplacian $B$ on $\mathcal{X}_{\sigma}^{\prime}$)
@@ 3. Solutions with $\varepsilon\Delta$ and Regularity
Here, $H^{2}(A)$ denotes the Sobolev space associated with the Neumann Stokes operator $A$, and $H^{4}(\Omega)$ is the standard Sobolev space equipped with the domain $\Omega$ (without boundary conditions).
@@
We start from the governing equation, the Neumann Stokes equation, written in terms of the operator $A$:
@@
Next, we estimate $\|A\partial_{t}u\|_{L^{2}(t_{0},T;L^{2})}$. By applying the operator $A$ and taking the $L^{2}(t_{0},T;L^{2}))$ norm on the equation for $\partial_{t}u$, we obtain the estimate:

### 20 Apr 2026 v2

@@ Abstract.
We consider the 3D incompressible Euler equations in bounded domains $\Omega$ with smooth boundary $\partial\Omega$. Based on [IMT-2019], we define the Besov space $B^{s}_{p,q}(A)$ by means of the Stokes operator AA with the Neumann boundary condition on $\partial\Omega$, and prove unique local existence theorem of strong solution for the initial data in the critical Besov space $B^{\frac{5}{2}}_{2,1}(A)$. Our proof relies on the method of vanishing viscosity. The commutator estimate plays an essential role for derivation of energy bounds which hold uniformly with respect to viscosity constants.
@@ 1. Introduction
In the present paper, we prove the local well-posedness in the critical Besov space $B^{\frac{5}{2}}_{2,1}(\Omega)$. Based on [IwKo-preprint] and [IMT-2019], we first define the Besov space $B^{s}_{p,q}(A)$ in terms of the Stokes operator AA in $L^{2}_{\sigma}(\Omega)$ with the Neumann boundary condition on $\partial\Omega$. Our method relies on the invisid limit ${\lim_{\varepsilon\to+0}u_{\varepsilon}}$ of a family $\{u_{\varepsilon}\}_{\varepsilon>0}$ of solutions to the following Navier-Stokes equations;
@@
It is essential to establish uniform bounds of $\{u_{\varepsilon}\}_{\varepsilon>0}$ in the Chemin-Lernar space $\tilde{L}^{\infty}(0,T;B^{\frac{5}{2}}_{2,1}(A))$. The commutator type estimate such as Kato-Ponce [KaPo-1988] plays a crucial role for such bounds. Since we make use of the Stokes operator AA with the Neumann boundary condition like (1.2), the boundary layer does not appear in the invisid limit.
@@
$\left\|u-\sum_{j\leq J}\phi_{j}(\sqrt{A})u\right\|_{L^{2}}\leq C2^{-\frac{5}{2}J}$
@@
This paper is organized as follows. In Section 2, we provide several preliminary lemmas and propositions required to prove Theorem $1.$ Section 3 is devoted to constructing solutions for the equation regularized by $\varepsilon\Delta$. In Section 4, we establish the uniform boundedness of the solutions obtained in Section 3 with respect to the parameter $\varepsilon>0$. Finally, in Section 5, we provide the proof of Theorem 1.1 by taking the limit as $\varepsilon\to 0$. Throughout the following sections, we suppose that $A_{2}$ is strictly positive.
@@ 2. Preliminary
In this section, we collect several preliminary definitions and results. In Subsection 2.1, we fundamental properties of the Stokes operator AA and the associated Besov spaces $B^{s}_{p,q}(A)$. Subsection 2.2 is devoted to providing several propositions and lemmas required for the proof of Theorem 1.1, including Sobolev-type embeddings, derivative estimates, and various properties of derivatives involving spectral restriction operators.
@@ 2.1. The Stokes operator and Besov spaces
We introduce the projection $\mathbb{P}$ onto the space of divergence-free vector fields and the Stokes operator on $\mathcal{X}_{\sigma}^{\prime}$. While the operator AA and its resolvent act on divergence-free fields, $B$ and its resolvent act on a more general class of vector fields.
Definition. (The projection $\mathbb{P}$, the Stokes operator AA and the Laplacian $B$ on $\mathcal{X}_{\sigma}^{\prime}$)
@@ 3. Solutions with $\varepsilon\Delta$ and Regularity
Here, $H^{2}(A)$ denotes the Sobolev space associated with the Neumann Stokes operator AA, and $H^{4}(\Omega)$ is the standard Sobolev space equipped with the domain $\Omega$ (without boundary conditions).
@@
We start from the governing equation, the Neumann Stokes equation, written in terms of the operator AA:
@@
Next, we estimate $\|A\partial_{t}u\|_{L^{2}(t_{0},T;L^{2})}$. By applying the operator AA and taking the $L^{2}(t_{0},T;L^{2}))$ norm on the equation for $\partial_{t}u$, we obtain the estimate: